OpenStudy (anonymous):

Find the standard form or the equation of the ellipse with vertices (±7, 0) and eccentricity e=2/7

OpenStudy (dumbcow):

 so a=7 e = c/a = 2/7 solve for c --> c/7 = 2/7 --> c = 2 now find b^2 where b^2 +c^2 = a^2 b^2 + 4 = 49 b^2 = 45 ellipse equation: \[\frac{x^{2}}{a^{2}}+\frac{y^{2}}{b^{2}}=1\]